# envname-P envname#1

### v1

@@ 5.4 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 5.3), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 3.2. We obtained 154,818 functions, of which 57,746 samples include docstrings. From this set, we utilized 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small-context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa hu2021lora and used 10% of the training data as the validation set to select the best checkpoint.
@@ 7 Limitations
Several code generation benchmarks utilize match-based metrics like Edit similarity (ES), BLEU, and CodeBLEU for evaluation (ding-etal-2024-cocomic-code; DBLP:conf/icml/ShrivastavaLT23; DBLP:conf/nips/DingWADTJRNBRX23; liao2023context; yin2018learning; iyer2018mapping). These metrics are straightforward to apply and may exhibit a strong correlation with execution metrics such as Pass@k. However, they cannot accurately measure functional correctness. For instance, comparing two Python code snippets where the only difference is the ":" character could result in a high ES and BLEU score. Nevertheless, one snippet may contain a syntax error, highlighting a limitation in these metrics for assessing true functionality.
@@ Appendix B Coverage Enhancement Effectiveness
Weak unit tests may inadvertently allow incorrect implementations to be determined as correct. Even with human-written tests, the overlooked coverage rates lead to evaluations that are incomplete and potentially misleading liu2023is. We present evidence supporting this argument, underscoring the limitations of prior work on code generation within repository-level contexts. As depicted in Figure 4, enhancing the number of test cases and coverage rates leads to a significant increase in the identification of incorrect generated solutions, causing the Pass@1 score to drop markedly (by over 5%). We investigated several solutions and found that most of the generated results did not fully utilize the given context (considered as human-provided). Instead, they primarily focused on addressing the problem described in the given natural language description. This indirectly overlooks the developer’s intentions, such as testing edge or corner cases, highlighting the limitations in following and understanding the provided intent and dependency context in these models. In summary, these findings underscore the effectiveness and importance of maintaining high-quality test cases for evaluation purposes.
@@ D.1 BasePrompt
Example of Full Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): """ ␣␣␣␣Customerrorraisedwhenreceivedobjectisnotastringasexpected. ␣␣␣␣""" def __init__(self, input_data: Any): """ ␣␣␣␣␣␣␣␣:paraminput_data:Anyreceivedobject ␣␣␣␣␣␣␣␣""" type_name = type(input_data).__name__ msg = ’Expected"str",received"{}"’.format(type_name) super().__init__(msg) def is_string(obj: Any) -> bool: """ ␣␣␣␣Checksifanobjectisastring. ␣␣␣␣*Example:* ␣␣␣␣>>>is_string(’foo’)#returnstrue ␣␣␣␣>>>is_string(b’foo’)#returnsfalse ␣␣␣␣:paramobj:Objecttotest. ␣␣␣␣:return:Trueifstring,falseotherwise. ␣␣␣␣""" return isinstance(obj, str) def is_camel_case(input_string: Any) -> bool: """ ␣␣␣␣Checksifastringisformattedascamelcase. ␣␣␣␣Astringisconsideredcamelcasewhen: ␣␣␣␣-it’scomposedonlybyletters([a-zA-Z])andoptionallynumbers([0-9]) ␣␣␣␣-itcontainsbothlowercaseanduppercaseletters ␣␣␣␣-itdoesnotstartwithanumber ␣␣␣␣*Examples:* ␣␣␣␣>>>is_camel_case(’MyString’)#returnstrue ␣␣␣␣>>>is_camel_case(’mystring’)#returnsfalse ␣␣␣␣:paraminput_string:Stringtotest. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Trueforacamelcasestring,falseotherwise. ␣␣␣␣""" return is_full_string(input_string) and CAMEL_CASE_TEST_RE.match(input_string) is not None def camel_case_to_snake(input_string, separator=’_’): """ ␣␣␣␣Convertacamelcasestringintoasnakecaseone. ␣␣␣␣(Theoriginalstringisreturnedifisnotavalidcamelcasestring) ␣␣␣␣*Example:* ␣␣␣␣>>>camel_case_to_snake(’ThisIsACamelStringTest’)#returns’this_is_a_camel_case_string_test’ ␣␣␣␣:paraminput_string:Stringtoconvert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:paramseparator:Signtouseasseparator. ␣␣␣␣:typeseparator:str ␣␣␣␣:return:Convertedstring. ␣␣␣␣"""
Example of Medium Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): """ ␣␣␣␣Customerrorraisedwhenreceivedobjectisnotastringasexpected. ␣␣␣␣""" def __init__(self, input_data: Any): """ ␣␣␣␣␣␣␣␣:paraminput_data:Anyreceivedobject ␣␣␣␣␣␣␣␣""" def is_string(obj: Any) -> bool: """ ␣␣␣␣Checksifanobjectisastring. ␣␣␣␣*Example:* ␣␣␣␣$>$is_string(’foo’)#returnstrue ␣␣␣␣$>$is_string(b’foo’)#returnsfalse ␣␣␣␣:paramobj:Objecttotest. ␣␣␣␣:return:Trueifstring,falseotherwise. ␣␣␣␣""" def is_camel_case(input_string: Any) -> bool: """ ␣␣␣␣Checksifastringisformattedascamelcase. ␣␣␣␣Astringisconsideredcamelcasewhen: ␣␣␣␣-it’scomposedonlybyletters([a-zA-Z])andoptionallynumbers([0-9]) ␣␣␣␣-itcontainsbothlowercaseanduppercaseletters ␣␣␣␣-itdoesnotstartwithanumber ␣␣␣␣*Examples:* ␣␣␣␣$>$is_camel_case(’MyString’)#returnstrue ␣␣␣␣$>$is_camel_case(’mystring’)#returnsfalse ␣␣␣␣:paraminput_string:Stringtotest. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Trueforacamelcasestring,falseotherwise. ␣␣␣␣""" def camel_case_to_snake(input_string, separator=’_’): """ ␣␣␣␣Convertacamelcasestringintoasnakecaseone. ␣␣␣␣(Theoriginalstringisreturnedifisnotavalidcamelcasestring) ␣␣␣␣*Example:* ␣␣␣␣$>$camel_case_to_snake(’ThisIsACamelStringTest’)#returns’this_is_a_camel_case_string_test’ ␣␣␣␣:paraminput_string:Stringtoconvert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:paramseparator:Signtouseasseparator. ␣␣␣␣:typeseparator:str ␣␣␣␣:return:Convertedstring. ␣␣␣␣"""
@@
Example of Short Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ ␣␣␣␣Convertacamelcasestringintoasnakecaseone. ␣␣␣␣(Theoriginalstringisreturnedifisnotavalidcamelcasestring) ␣␣␣␣*Example:* ␣␣␣␣$>$camel_case_to_snake(’ThisIsACamelStringTest’)#returns’this_is_a_camel_case_string_test’ ␣␣␣␣:paraminput_string:Stringtoconvert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:paramseparator:Signtouseasseparator. ␣␣␣␣:typeseparator:str ␣␣␣␣:return:Convertedstring. ␣␣␣␣"""
@@ Appendix E Studied LLMs: Supplemental results
Indeed, Figure 5 illustrates the proportion of generated functions that are empty for each LLM using BasePrompt. The findings indicate that utilizing Medium context results in a substantial number of empty functions. This may be due to the input format of the context when using BasePrompt, which can mislead the model into interpreting dependency functions as few-shot examples. In the Medium context, the function bodies of dependencies are removed, making their format identical to the target function prompt. This similarity can mislead the LMs, resulting in empty solutions. Particularly, Starcoder-2 is heavily impacted by this issue, as over 31% of its generated results are empty functions, revealing a significant weakness of the model. Meanwhile, short context effectively decreases the occurrence of empty function generation by the model and, in certain instances, improves models’ ability in dependency calls (e.g. CodeLlama-13b-Python, WizardCoder-Python-13B-V1.0, and Mixtral-8x7B-Instruct-v0.1 in Table 3). We believe that the following reasons could contribute to this observation. Firstly, employing short context reduces the input token count, preventing truncation when exceeding the maximum length limit, thus allowing uninterrupted solution generation by the model. This reduced context enables models to concentrate exclusively on dependency signatures, thereby enhancing the probability that generated solutions effectively utilize these dependency token names. Moreover, function names hold substantial semantic value by delineating the function’s purpose. Many studies have underscored that code summarization heavily relies on extracting information from function names (haldar2024analyzing; mondal-etal-2023-robust; sontakke2022code). Therefore, this concise representation of dependencies has the potential to improve how models utilize dependencies in generating code.
@@ Appendix F Multi-round Debugging
# Prompt: import base64 import random import unicodedata ... class InvalidInputError(TypeError): ... def is_string(obj: Any) -> bool: ... def reverse(input_string: str) -> str: """ ␣␣␣␣Returnsthestringwithitscharsreversed. ␣␣␣␣*Example:* ␣␣␣␣$>$reverse(’hello’)#returns’olleh’ ␣␣␣␣:paraminput_string:Stringtorevert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Reversedstring. ␣␣␣␣""" ------------------------------------------------------- # Solution: def reverse(input_string: str) -> str: """ ␣␣␣␣Returnsthestringwithitscharsreversed. ␣␣␣␣*Example:* ␣␣␣␣$>$reverse(’hello’)#returns’olleh’ ␣␣␣␣:paraminput_string:Stringtorevert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Reversedstring. ␣␣␣␣""" if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1] # Model type: Pretrained # Passed: No def reverse(input_string: str) -> str: """ ␣␣␣␣Returnsthestringwithitscharsreversed. ␣␣␣␣*Example:* ␣␣␣␣$>$reverse(’hello’)#returns’olleh’ ␣␣␣␣:paraminput_string:Stringtorevert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Reversedstring. ␣␣␣␣""" return input_string[::-1] ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def reverse(input_string: str) -> str: """ ␣␣␣␣Returnsthestringwithitscharsreversed. ␣␣␣␣*Example:* ␣␣␣␣$>$reverse(’hello’)#returns’olleh’ ␣␣␣␣:paraminput_string:Stringtorevert. ␣␣␣␣:typeinput_string:str ␣␣␣␣:return:Reversedstring. ␣␣␣␣""" if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1]
@@
# Prompt: import random import luigi import luigi.configuration ... class hadoopcli(luigi.Config): command = ... version = ... def get_configured_hadoop_version(): """ ␣␣␣␣CDH4(hadoop2+)hasaslightlydifferentsyntaxforinteractingwithhdfs ␣␣␣␣viathecommandline. ␣␣␣␣ThedefaultversionisCDH4,butonecanoverride ␣␣␣␣thissettingwith"cdh3"or"apache1"inthehadoopsectionoftheconfig ␣␣␣␣inordertousetheoldsyntax. ␣␣␣␣""" ------------------------------------------------------- # Solution: def get_configured_hadoop_version(): """ ␣␣␣␣CDH4(hadoop2+)hasaslightlydifferentsyntaxforinteractingwithhdfs ␣␣␣␣viathecommandline. ␣␣␣␣ThedefaultversionisCDH4,butonecanoverride ␣␣␣␣thissettingwith"cdh3"or"apache1"inthehadoopsectionoftheconfig ␣␣␣␣inordertousetheoldsyntax. ␣␣␣␣""" return hadoopcli().version.lower() # Model type: Pretrained # Passed: Yes def get_configured_hdfs_client(): """ ␣␣␣␣Thisisahelperthatfetchestheconfigurationvaluefor’client’in ␣␣␣␣the[hdfs]section.Itwillreturntheclientthatretainsbackwards ␣␣␣␣compatibilitywhen’client’isn’tconfigured. ␣␣␣␣""" return luigi.configuration.get_config().get(’hdfs’, ’client’, ’hadoopcli’) ------------------------------------------------------- # Model type: Instruction-tuned # Passed: No def get_configured_hadoop_version(): """ ␣␣␣␣CDH4(hadoop2+)hasaslightlydifferentsyntaxforinteractingwithhdfs ␣␣␣␣viathecommandline. ␣␣␣␣ThedefaultversionisCDH4,butonecanoverride ␣␣␣␣thissettingwith"cdh3"or"apache1"inthehadoopsectionoftheconfig ␣␣␣␣inordertousetheoldsyntax. ␣␣␣␣""" hadoop_cli = hadoopcli().command.split() hadoop_version = hadoopcli().version if hadoop_version == "cdh3": return hadoop_cli + ["fs"] elif hadoop_version == "apache1": return hadoop_cli + ["dfs"] else: return hadoop_cli
@@
# Prompt: import asyncio import builtins import collections ... def is_future(x: Any) -> bool: return isinstance(x, FUTURES) def _create_future() -> Future: future = Future() # type: Future ... def maybe_future(x: Any) -> Future: """Converts‘‘x‘‘intoa‘.Future‘. ␣␣␣␣If‘‘x‘‘isalreadya‘.Future‘,itissimplyreturned;otherwise ␣␣␣␣itiswrappedinanew‘.Future‘.␣␣Thisissuitableforuseas ␣␣␣␣‘‘result=yieldgen.maybe_future(f())‘‘whenyoudon’tknowwhether ␣␣␣␣‘‘f()‘‘returnsa‘.Future‘ornot. ␣␣␣␣..deprecated::4.3 ␣␣␣␣Thisfunctiononlyhandles‘‘Futures‘‘,nototheryieldableobjects. ␣␣␣␣Insteadof‘maybe_future‘,checkforthenon-futureresulttypes ␣␣␣␣youexpect(oftenjust‘‘None‘‘),and‘‘yield‘‘anythingunknown. ␣␣␣␣""" ------------------------------------------------------- # Solution: def maybe_future(x: Any) -> Future: """Converts‘‘x‘‘intoa‘.Future‘. ␣␣␣␣If‘‘x‘‘isalreadya‘.Future‘,itissimplyreturned;otherwise ␣␣␣␣itiswrappedinanew‘.Future‘.␣␣Thisissuitableforuseas ␣␣␣␣‘‘result=yieldgen.maybe_future(f())‘‘whenyoudon’tknowwhether ␣␣␣␣‘‘f()‘‘returnsa‘.Future‘ornot. ␣␣␣␣..deprecated::4.3 ␣␣␣␣␣␣␣Thisfunctiononlyhandles‘‘Futures‘‘,nototheryieldableobjects. ␣␣␣␣␣␣␣Insteadof‘maybe_future‘,checkforthenon-futureresulttypes ␣␣␣␣␣␣␣youexpect(oftenjust‘‘None‘‘),and‘‘yield‘‘anythingunknown. ␣␣␣␣""" if is_future(x): return x else: fut = _create_future() fut.set_result(x) return fut # Model type: Pretrained # Passed: Yes def maybe_future(x: Any) -> Future: """Converts‘‘x‘‘intoa‘.Future‘. ␣␣␣␣If‘‘x‘‘isalreadya‘.Future‘,itissimplyreturned;otherwise ␣␣␣␣itiswrappedinanew‘.Future‘.␣␣Thisissuitableforuseas ␣␣␣␣‘‘result=yieldgen.maybe_future(f())‘‘whenyoudon’tknowwhether ␣␣␣␣‘‘f()‘‘returnsa‘.Future‘ornot. ␣␣␣␣..deprecated::4.3 ␣␣␣␣Thisfunctiononlyhandles‘‘Futures‘‘,nototheryieldableobjects. ␣␣␣␣Insteadof‘maybe_future‘,checkforthenon-futureresulttypes ␣␣␣␣youexpect(oftenjust‘‘None‘‘),and‘‘yield‘‘anythingunknown. ␣␣␣␣""" if is_future(x): return x future = Future() # type: Future future.set_result(x) return future ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def maybe_future(x: Any) -> Future: """Converts‘‘x‘‘intoa‘.Future‘. ␣␣␣␣If‘‘x‘‘isalreadya‘.Future‘,itissimplyreturned;otherwise ␣␣␣␣itiswrappedinanew‘.Future‘.␣␣Thisissuitableforuseas ␣␣␣␣‘‘result=yieldgen.maybe_future(f())‘‘whenyoudon’tknowwhether ␣␣␣␣‘‘f()‘‘returnsa‘.Future‘ornot. ␣␣␣␣..deprecated::4.3 ␣␣␣␣Thisfunctiononlyhandles‘‘Futures‘‘,nototheryieldableobjects. ␣␣␣␣Insteadof‘maybe_future‘,checkforthenon-futureresulttypes ␣␣␣␣youexpect(oftenjust‘‘None‘‘),and‘‘yield‘‘anythingunknown. ␣␣␣␣""" if isinstance(x, Future): return x elif isawaitable(x): return asyncio.ensure_future(x) else: future = Future() future.set_result(x) return future

### v2

@@ 5.4 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 5.3), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 3.2. We obtained 154,818 functions, of which 57,746 samples include docstrings. From this set, we utilized 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small-context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa Hu et al. (2021) and used 10% of the training data as the validation set to select the best checkpoint.
@@ 7 Limitations
Several code generation benchmarks utilize match-based metrics like Edit similarity (ES), BLEU, and CodeBLEU for evaluation [Ding et al., 2024, Shrivastava et al., 2023, Ding et al., 2023, Liao et al., 2023, Yin et al., 2018, Iyer et al., 2018]. These metrics are straightforward to apply and may exhibit a strong correlation with execution metrics such as Pass@k. However, they cannot accurately measure functional correctness. For instance, comparing two Python code snippets where the only difference is the ":" character could result in a high ES and BLEU score. Nevertheless, one snippet may contain a syntax error, highlighting a limitation in these metrics for assessing true functionality.
@@ Appendix B Coverage Enhancement Effectiveness
Weak unit tests may inadvertently allow incorrect implementations to be determined as correct. Even with human-written tests, the overlooked coverage rates lead to evaluations that are incomplete and potentially misleading Liu et al. [2023a]. We present evidence supporting this argument, underscoring the limitations of prior work on code generation within repository-level contexts. As depicted in Figure 5, enhancing the number of test cases and coverage rates leads to a significant increase in the identification of incorrect generated solutions, causing the Pass@1 score to drop markedly (by over 5%). We investigated several solutions and found that most of the generated results did not fully utilize the given context (considered as human-provided). Instead, they primarily focused on addressing the problem described in the given natural language description. This indirectly overlooks the developer’s intentions, such as testing edge or corner cases, highlighting the limitations in following and understanding the provided intent and dependency context in these models. In summary, these findings underscore the effectiveness and importance of maintaining high-quality test cases for evaluation purposes.
@@ D.1 BasePrompt
Example of Full Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): """ Custom error raised when received object is not a string as expected. """ def __init__(self, input_data: Any): """ :param input_data: Any received object """ type_name = type(input_data).__name__ msg = ’Expected "str", received "{}"’.format(type_name) super().__init__(msg) def is_string(obj: Any) -> bool: """ Checks if an object is a string. *Example:* >>> is_string(’foo’) # returns true >>> is_string(b’foo’) # returns false :param obj: Object to test. :return: True if string, false otherwise. """ return isinstance(obj, str) def is_camel_case(input_string: Any) -> bool: """ Checks if a string is formatted as camel case. A string is considered camel case when: - it’s composed only by letters ([a-zA-Z]) and optionally numbers ([0-9]) - it contains both lowercase and uppercase letters - it does not start with a number *Examples:* >>> is_camel_case(’MyString’) # returns true >>> is_camel_case(’mystring’) # returns false :param input_string: String to test. :type input_string: str :return: True for a camel case string, false otherwise. """ return is_full_string(input_string) and CAMEL_CASE_TEST_RE.match(input_string) is not None def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* >>> camel_case_to_snake(’ThisIsACamelStringTest’) # returns ’this_is_a_camel_case_string_test’ :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@
Example of Medium Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): """ Custom error raised when received object is not a string as expected. """ def __init__(self, input_data: Any): """ :param input_data: Any received object """ def is_string(obj: Any) -> bool: """ Checks if an object is a string. *Example:* $>$ is_string(’foo’) # returns true $>$ is_string(b’foo’) # returns false :param obj: Object to test. :return: True if string, false otherwise. """ def is_camel_case(input_string: Any) -> bool: """ Checks if a string is formatted as camel case. A string is considered camel case when: - it’s composed only by letters ([a-zA-Z]) and optionally numbers ([0-9]) - it contains both lowercase and uppercase letters - it does not start with a number *Examples:* $>$ is_camel_case(’MyString’) # returns true $>$ is_camel_case(’mystring’) # returns false :param input_string: String to test. :type input_string: str :return: True for a camel case string, false otherwise. """ def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns ’this_is_a_camel_case_string_test’ :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
Example of Short Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns ’this_is_a_camel_case_string_test’ :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@ Appendix E Studied LLMs: Supplemental results
Indeed, Figure 6 illustrates the proportion of generated functions that are empty for each LLM using BasePrompt. The findings indicate that utilizing Medium context results in a substantial number of empty functions. This may be due to the input format of the context when using BasePrompt, which can mislead the model into interpreting dependency functions as few-shot examples. In the Medium context, the function bodies of dependencies are removed, making their format identical to the target function prompt. This similarity can mislead the LMs, resulting in empty solutions. Particularly, Starcoder-2 is heavily impacted by this issue, as over 31% of its generated results are empty functions, revealing a significant weakness of the model. Meanwhile, short context effectively decreases the occurrence of empty function generation by the model and, in certain instances, improves models’ ability in dependency calls (e.g. CodeLlama-13b-Python, WizardCoder-Python-13B-V1.0, and Mixtral-8x7B-Instruct-v0.1 in Table 3). We believe that the following reasons could contribute to this observation. Firstly, employing short context reduces the input token count, preventing truncation when exceeding the maximum length limit, thus allowing uninterrupted solution generation by the model. This reduced context enables models to concentrate exclusively on dependency signatures, thereby enhancing the probability that generated solutions effectively utilize these dependency token names. Moreover, function names hold substantial semantic value by delineating the function’s purpose. Many studies have underscored that code summarization heavily relies on extracting information from function names [Haldar and Hockenmaier, 2024, Mondal et al., 2023, Sontakke et al., 2022]. Therefore, this concise representation of dependencies has the potential to improve how models utilize dependencies in generating code.
@@ Appendix F Multi-round Debugging
# Prompt: import base64 import random import unicodedata ... class InvalidInputError(TypeError): ... def is_string(obj: Any) -> bool: ... def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ ------------------------------------------------------- # Solution: def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1] # Model type: Pretrained # Passed: No def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ return input_string[::-1] ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1]
@@
# Prompt: import random import luigi import luigi.configuration ... class hadoopcli(luigi.Config): command = ... version = ... def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ ------------------------------------------------------- # Solution: def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ return hadoopcli().version.lower() # Model type: Pretrained # Passed: Yes def get_configured_hdfs_client(): """ This is a helper that fetches the configuration value for ’client’ in the [hdfs] section. It will return the client that retains backwards compatibility when ’client’ isn’t configured. """ return luigi.configuration.get_config().get(’hdfs’, ’client’, ’hadoopcli’) ------------------------------------------------------- # Model type: Instruction-tuned # Passed: No def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ hadoop_cli = hadoopcli().command.split() hadoop_version = hadoopcli().version if hadoop_version == "cdh3": return hadoop_cli + ["fs"] elif hadoop_version == "apache1": return hadoop_cli + ["dfs"] else: return hadoop_cli
@@
# Prompt: import asyncio import builtins import collections ... def is_future(x: Any) -> bool: return isinstance(x, FUTURES) def _create_future() -> Future: future = Future() # type: Future ... def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ ------------------------------------------------------- # Solution: def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x else: fut = _create_future() fut.set_result(x) return fut # Model type: Pretrained # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x future = Future() # type: Future future.set_result(x) return future ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if isinstance(x, Future): return x elif isawaitable(x): return asyncio.ensure_future(x) else: future = Future() future.set_result(x) return future